OpenStudy (anonymous):

Input the equation of the given line in standard form. The line including the two points (5, 3), (8, -2).? Help on how to find the answer?

OpenStudy (texaschic101):

first find the slope using the points given and the slope formula slope(m) = (y2 - y1) / (x2 - x1) (5,3)...x1 = 5 and y1 = 3 (8,-2)...x2 = 8 and y2 = -2 now we sub slope(m) = (-2 - 3) / (8 - 5) slope(m) = -5/3 now that we know the slope, we need the y intercept...that can be found by using the formula y = mx + b and either of your sets of points. y = mx + b slope(m) = -5/3 (5,3)...x = 5 and y = 3 now we sub and solve for b, your y intercept 3 = -5/3(5) + b 3 = -25/3 + b 3 + 25/3 = b -- common denominator is 3 9/3 + 25/3 = b 34/3 = b so our equation is : y = -5/3 + 34/3 However, we need it in standard form Ax + By = C y = -5/3x + 34/3 -- multiply by the common denominator 3 3y = -5x + 34 -- add 5x to both sides 5x + 3y = 34 <== standard form